### v1

# Re-Introducing LayerNorm: Geometric Meaning, Irreversibility and a Comparative Study with RMSNorm

Abstract
Layer normalization is a pivotal step in the transformer architecture. This paper delves into the less explored geometric implications of this process, examining how LayerNorm influences the norm and orientation of hidden vectors in the representation space. We show that the definition of LayerNorm is innately linked to the uniform vector, defined as $\bm{1}=[1,1,1,1,\cdots,1]^{T}\in\mathbb{R}^{d}$. We then show that the standardization step in LayerNorm can be understood in three simple steps: (i) remove the component of a vector along the uniform vector, (ii) normalize the remaining vector, and (iii) scale the resultant vector by $\sqrt{d}$, where $d$ is the dimensionality of the representation space. We also introduce the property of ”irreversibility” for LayerNorm, where we show that the information lost during the normalization process cannot be recovered. In other words, unlike batch normalization, LayerNorm cannot learn an identity transform. While we present possible arguments for removing the component along the uniform vector, the choice of removing this component seems arbitrary and not well motivated by the original authors. To evaluate the usefulness of this step, we compare the hidden representations of LayerNorm-based LLMs with models trained using RMSNorm and show that all LLMs naturally align representations orthogonal to the uniform vector, presenting the first mechanistic evidence that removing the component along the uniform vector in LayerNorm is a redundant step. Our findings support the use of RMSNorm over LayerNorm as it is not only more computationally efficient with comparable downstream performance, but also learns a similar distribution of hidden representations that operate orthogonal to the uniform vector.
1 Introduction
The transformer architecture (Vaswani et al., 2017) has been the cornerstone of most recent advances in artificial intelligence and has rapidly become the architecture of choice in natural language processing, computer vision, speech, and various other domains. Layer normalization, a crucial yet often overlooked component of the transformer architecture, plays an integral role in stabilizing the training process in transformer-based models. Introduced by Ba et al. (2016), layer normalization standardizes the features at each layer, adjusting and scaling the activations to have zero mean and unit variance within a vector. This normalization is performed independently for each hidden vector in contrast to batch normalization (Ioffe & Szegedy, 2015), which relies on statistics (mean and variance) from a batch of data points. Layer normalization is especially effective when working with long sequences of variable lengths. While the benefits of layer normalization, such as improved training speed and better convergence, are well documented (Ba et al., 2016; Xu et al., 2019; Zhang & Sennrich, 2019), its specific effects on the hidden vectors within a model and the global properties of the resulting representations remain surprisingly underexplored.
In this paper, we first discuss the global effects of layer normalization on a vector. The conventional explanation of the LayerNorm operation is usually as follows: standardize each vector by subtracting the mean of its elements, divide by the standard deviation. While this is an accurate procedural definition, we ask a more global question: How does the LayerNorm operation transform a vector in representation space? We present a novel interpretation of LayerNorm and show that it can be understood in three steps: (i) throw away the component of the given vector along $\bm{1}=[1,1,1,1,\cdots,1]^{T}\in\mathbb{R}^{d}$, (ii) normalize the resultant vector; and (iii) scale the resulting vector by $\sqrt{d}$, where $d$ is the dimensionality of the vector space. This process is illustrated in Figure 1. This operation throws away the information along $\bm{1}=[1,1,1,1,\cdots 1]^{T}$, which we call the uniform vector, indicating that the information along the uniform vector may not be important.
We then present the property of “irreversibility,” where we show that the information lost in the LayerNorm process cannot be recovered. While “reversibility” is an established property in BatchNorm (Ioffe & Szegedy, 2015), this property has not been discussed in literature for layer normalization to the best of our knowledge. These results naturally lead to discussion about the importance of information along the uniform vector, which is lost during LayerNorm irreversibly, and its comparison with RMSNorm (Zhang & Sennrich, 2019), a variant of LayerNorm that doesn’t remove the component along the uniform vector and is used to train the latest Llama models (Touvron et al., 2023a; b; Dubey et al., 2024). Figure 2(a) shows the difference between the two normalization methods. We empirically show that RMSNorm-based models naturally produce hidden representations orthogonal to the uniform vector, thus rendering the removal of the uniform vector in LayerNorm inconsequential in practice.
(a) This figure shows the original vector (in blue) and the uniform vector (in red) in a 3-D space.
(b) This figure shows the projection of the original vector along the uniform vector, $(\bm{x}.\hat{\bm{1}})\hat{\bm{1}}$, and the remaining component, $\bm{x^{\bm{1}}_{\perp}}$.
(c) The component of the original vector, after removing the projection along the uniform vector, is kept ( $\bm{x^{\bm{1}}_{\perp}}$, shown in yellow).
(d) Normalizing perpendicular component to the uniform vector ($\bm{x^{\bm{1}}_{\perp}}$) to unit norm.
(e) Scale of resultant vector by $\sqrt{d}$, where $d$ is the dimensionality of the representation space.
(f) Finally, the scale-and-shift step, which scales and shifts the resulting vector according to learnt parameters.
Figure 1: Visualization of LayerNorm operation on a random original vector
To summarize, in this paper, we make the following contributions:
  1. 1.
    Provide a simple geometric interpretation of the layer normalization process that provides an intuitive explanation of how LayerNorm transforms a vector in representation space.
  2. 2.
    Introduce the property of “irreversibility” for LayerNorm and show that information lost in the layer normalization process cannot be recovered.
  3. 3.
    Show that removing the component along the uniform vector is a redundant step in LayerNorm, and models trained using RMSNorm also learn to create representations orthogonal to the uniform vector. These findings support using RMSNorm over LayerNorm.
  4. 4.
    Finally, we also present LayerNorm in action with experiments depicting its importance in stabilizing the residual stream vectors. LayerNorm not only regulates the ever-increasing norm of hidden vectors inside LLMs but also rotates each vector to orient it properly within representation space.
2 Background
Let $\bm{x}\in\mathbb{R}^{d}$ represent an intermediate hidden representation in a transformer-based model, on which the LayerNorm operation is applied. Then the following steps summarize the layer normalization operation:
  1. 1.
    Calculate the mean, $\mu=\frac{1}{d}\sum^{d}_{i=1}x_{i}$, and standard deviation, $\sigma=\sqrt{\frac{1}{d}\sum^{d}_{i=1}(x_{i}-\mu)^{2}}$, where $x_{i}$ are the components of $\bm{x}$
  2. 2.
    Standardize the components of the hidden vector $\bm{x}$ to get $\bm{y}$, such that $y_{i}=\frac{x_{i}-\mu}{\sigma}$, $\forall x_{i}\in\bm{x}$
  3. 3.
    Scale and shift to get $\bm{z}=\bm{\alpha}\odot\bm{y}+\bm{\beta}$, where $\odot$ represents element-wise multiplication of two vectors, and $\bm{\alpha}$, $\bm{\beta}$ are scaling and shifting vector parameters learned during training.
In this paper, we refer to the combination of step-1 and step-2 as the ”standardization step”, whereas step-3 is referred to as the ”scale-and-shift” step.
2.1 Computations in Decoder-only LLMs
In this paper, we study the hidden representations of modern decoder-only large language models (LLMs). Let $h^{l}$ represent the intermediate hidden state vectors between each decoder layer. As depicted in Figure 2(b), the computations within a layer of most decoder-only LLMs proceed as follows:
[…]
$a^{l}$ $=\texttt{Att}(f^{l})$ (2)
[…]
$m^{l}$ $=W^{l}_{proj}\sigma(W^{l}_{fc}g^{l}+b^{l}_{fc})+b_{proj}$ (4)
$h^{l}$ $=h^{l-1}+a^{l}+m^{l}$ (5)
(a) A diagrammatic explanation of LayerNorm and RMSNorm.
(b) A high-level diagram representing the computation within one decoder block of an LLM.
Figure 2: Combined figures showing the diagrams for LayerNorm and RMSNorm alongside the computation within a decoder block of an LLM.
The intermediate hidden vectors between each layer, $h^{l}$, are also sometimes called the residual stream. Let LN1 represent the first LayerNorm function that acts just before the attention module and LN2 represent the second LayerNorm just before the MLP module. We abstract out the computations of the attention and MLP modules to focus on the layer normalization blocks.
As the vectors computed in the attention and MLP modules get added to the residual stream at each layer, the residual stream represents a summation of an increasing number of vectors. A non-recursive formula for the residual stream depicts this clearly:
$h^{l}=h^{0}+\sum^{i=l}_{i=0}+a^{i}+\sum^{i=l}_{i=0}+m^{i}$ (6)
The residual stream thus represents a continuous summation of vectors computed at the attention and MLP modules in each layer with the incoming residual stream. This leads to an increasing norm of the residual stream (section 4.1), thus necessitating a normalization operation.
3 Re-Introducing Layer Normalization
While the original definition of LayerNorm (section 2) gives us an operational description, it tells us very little about the global properties of the resulting vectors. For example, with the steps described in section 2, we get very little insight about the norm or the orientation of the standardized vector ($\bm{y}$). All we know is that the components of the vector $\bm{y}$ are standardized to have zero mean and unit variance within the vector. But how is this standardized vector oriented compared to the original vector and what is its norms? Such an explanation was also absent in the original formulation of layer normalization (Ba et al., 2016).
Initial explorations into the geometry of layer normalization by Brody et al. (2023) demonstrated that $\bm{y}$ maintains a norm of $\sqrt{d}$ and is oriented orthogonal to the constant vector $\bm{1}=[1,1,1,1,\cdots,1]^{T}\in\mathbb{R}^{d}$, where $d$ represents the dimensionality of the vector space of hidden vectors. Furthermore, a popular variant of LayerNorm, known as RMSNorm (Zhang & Sennrich, 2019), which omits the “mean subtraction” in the standardization step, has come out as a viable alternative to LayerNorm. In this paper, we expand upon these foundational studies by presenting a novel interpretation of the global effects of LayerNorm on a vector, which we argue provides a more intuitive and informative description of how LayerNorm modifies a hidden vector in representation space.
To understand the global effect of LayerNorm on a vector $\bm{x}$, we need to represent the standardization steps in terms of the vector $\bm{x}$ and not its components. A neat way to write the mean of the components of a vector $\bm{x}$ is:
$\mu=\frac{1}{d}\bm{1}^{T}\bm{x}$ (7)
where $\bm{1}=[1,1,1,1,\cdots 1]^{T}$ such that $\bm{1}\in\mathbb{R}^{d}$, which we refer to as the uniform vector in this paper. Using the scalar mean value calculated above, we define the mean vector as $\bm{\mu}=\mu\bm{1}$, where $\mu$ is a scalar mean as calculated in equation 7. Thus, the mean vector is a vector with each component equal to the mean of the components of vector $\bm{x}$. Using this, we can also rewrite the standard deviation of the components of $\bm{x}$ as:
$\sigma=\sqrt{\frac{1}{d}(\bm{x}-\bm{\mu})^{T}(\bm{x}-\bm{\mu})}$ (8)
Once we do this, we can now write the standardize step in the LayerNorm algorithm (step-2) in vector form as follows:
$\bm{y}=\frac{1}{\sigma}(\bm{x}-\bm{\mu})$ (9)
In the above equation, we are assuming that the standard deviation is non-zero. As can be seen in equation 8, the standard deviation is zero only when $\bm{x}=\bm{\mu}$, which happens when all components of the hidden vector are equal. In practice, the authors of LayerNorm add an error term in the denominator to prevent this from happening. In the discussion that follows, we will assume that $\bm{x}\neq\bm{\mu}$, which is the same as saying that all components of $\bm{x}$ are not equal. This assumption does not lead to any loss of generality, as will be evident in later discussions. If the standard deviation is zero or $\bm{x}=\bm{\mu}$, then LayerNorm outputs a vector with all its components equal to zero.
3.1 The Uniform Vector and the Mean Vector
The vector definition of layer normalization in equation 9 requires us to define two new vectors: the uniform vector and the mean vector. Since these are non-standard vectors, it is important to understand the properties of these vectors.
The uniform vector or $\bm{1}=[1,1,1,1,\cdots 1]^{T}$, is called so because all its components are equal or uniform and set to 1. This vector plays a very important role in the formulation and understanding of layer normalization. An important thing to note here is the norm of the uniform vector: $\|\bm{1}\|_{2}=\sqrt{\bm{1}^{T}\bm{1}}=\sqrt{d}$.
The mean vector in the context of LayerNorm is defined in a non-traditional manner. Traditionally, a mean vector is the sum of a few vectors. But in our definition, the mean vector, $\bm{\mu}=\mu\bm{1}$, is a scaled version of the uniform vector. It is scaled by a value that is the mean or the average of all the components of the hidden vector and is oriented in the direction of the uniform vector. Note that the mean vector is a function of an underlying vector $\bm{x}$, as is implicit in the calculation of $\mu$ in equation 7. An interesting property of the mean vector is its norm. Let’s define $\theta_{x1}$ as the angle between vectors $\bm{x}$ and the uniform vector. Then:
$\begin{split}\|\bm{\mu}\|&=\|\mu\bm{1}\|=\mu\|\bm{1}\|\\ &=\Big{(}\frac{1}{d}\bm{1}^{T}\bm{x}\Big{)}\|\bm{1}\|\\ &=\frac{1}{d}\|\bm{1}\|^{2}\|\bm{x}\|\cos\theta_{x1}=\|\bm{x}\|\cos\theta_{x1}\end{split}$ (10)
Here we replace the formula for the mean from equation 7, expand the inner product between the uniform vector and $\bm{x}$ in terms of their norms and the angle between them, and use the fact that $\|\bm{1}\|_{2}=\sqrt{d}$.
The norm of the mean vector gives us very important insights about what’s going on in the layer normalization process. Equation 10 shows that the norm of the mean vector is nothing but the projection of the underlying vector $\bm{x}$ along the uniform vector. By definition, the mean vector, $\bm{\mu}=\mu\bm{1}$, is oriented along the direction of the uniform vector. In other words, given that $\theta$ is the angle between vectors $\bm{x}$ and the uniform vector,
$\begin{split}\|\bm{\mu}\|=\|\bm{x}\|\cos\theta_{x1}=\bm{x}.\hat{\bm{1}}\end{split}$ (11)
and
$\begin{split}\bm{\mu}\ =(\bm{x}.\hat{\bm{1}})\hat{\bm{1}}\end{split}$ (12)
To the best of our knowledge, the geometrical meaning of the mean vector has not been discussed in literature before our work. In the next section, we incorporate this information to explain the LayerNorm process, giving a very intuitive and informative description of the process.
3.2 Explanation of Layer Normalization
If we go through the steps of layer normalization as discussed in section 2, the component-wise subtraction of the mean of all components of a vector can now be written completely in terms of the vector $\bm{x}$. We define $\bm{x^{\bm{1}}_{\perp}}$ as the component of $\bm{x}$ orthogonal to the uniform vector as follows:
$\bm{x^{\bm{1}}_{\perp}}=\bm{x}-\bm{\mu}=\bm{x}-(\bm{x}.\hat{\bm{1}})\hat{\bm{1}}$ (13)
Thus, subtracting the mean of the components of a vector is the same as the removal of the projection of the vector along the uniform vector. The complete standardization step of the layer normalization operation can now be written as:
$\bm{y}=\sqrt{d}\frac{\bm{x^{\bm{1}}_{\perp}}}{\|\bm{x^{\bm{1}}_{\perp}}\|_{2}}$ (14)
where $d$ is the dimensionality of the vector being normalized. This shows that layer normalization can be simply defined as the normalization of the component of a vector orthogonal to the uniform vector, accompanied by a scaling factor. Figure 1 shows an intuitive description of this process. With this, we present the simple and elegant definition of the layer normalization process with a deep geometric meaning. The most intuitive recipe of the standardization step in LayerNorm is as follows:
The Standardization Step in LayerNorm
1:Throw away the component of a vector along the uniform vector, $\vec{\bm{1}}=[1,1,1,1,\cdots 1]^{T}$ (Fig 1(c))
2:Normalize the remaining vector (Fig 1(d))
3:Scale the resulting vector by $\sqrt{d}$, where $d$ is the dimensionality of the vector space (Fig 1(e))
Equation 14 also shows that the norm of the standardized vector is $\sqrt{d}$ and is oriented orthogonal to the direction of the uniform vector, which means it exists in a $d-1$ dimensional subspace orthogonal to the uniform vector (Brody et al., 2023).
3.3 The Irreversibility of Layer Normalization
The idea of reversibility is discussed briefly while introducing batch normalization (Ioffe & Szegedy, 2015), which normalizes each feature dimension in the input vector $\bm{x}$ independently by calculating the mean and standard deviation statistics over a large sample of such vectors. Specifically, let $\bm{x}^{1},\bm{x}^{2},\dots,\bm{x}^{b}$ be a set of $b$ vectors in the training set, each vector of dimension $d$. Then, both the batch normalization and layer normalization processes for each component can be represented in the following two-step process:
[…]
The key difference between batch normalization and layer normalization lies in how the expectation values and variance are calculated for each component. In the case of batch normalization, the mean and variance for each dimension are calculated over the training set, i.e.
$\mathbb{E}[x_{i}]$ $=\frac{1}{b}\sum^{j=b}_{j=1}x^{j}_{i}$ (17)
$Var(x_{i})$ $=\frac{1}{b}\sum^{j=b}_{j=1}(x^{j}_{i}-\mathbb{E}[x^{j}_{i}])^{2}$ (18)
Note that for batch normalization, $\mathbb{E}[x_{i}]$ and $Var(x_{i})$ are the same for each vector $\bm{x}^{j}\in\{\bm{x}^{1},\bm{x}^{2},\dots,\bm{x}^{b}\}$, but different for each component of the vector. In total, for a $d$-dimensional hidden vector space, there are $2d$ statistics in batch normalization, $d$ for the mean and $d$ for variance. This means that the information lost during the standardization step in batch normalization, if important, can be recovered in the scale-and-shift step by simply learning $\alpha_{i}=\sqrt{Var(x_{i})}$ and $\beta_{i}=\mathbb{E}[x_{i}]$, since we have $2d$ learnable parameters in batch normalization. The scale-and-shift step was a very conscious design choice of the inventors of batch normalization to allow for batch normalization to represent an identity transformation if the original hidden representations were optimal for the network.
A subtle but very important difference with layer normalization is how the expectation and variances are calculated for each component. For layer normalization, these statistics are calculated for each vector independently. Thus,
$\mathbb{E}[x_{i}]$ $=\mathbb{E}[x^{j}_{i}]=\frac{1}{d}\sum^{i=d}_{i=1}x^{j}_{i}$ (19)
$Var(x_{i})$ $=Var(x^{j}_{i})\frac{1}{d}\sum^{i=b}_{i=1}(x^{j}_{i}-\mathbb{E}[x^{j}_{i}])^{2}$ (20)
We recommend the reader carefully note the summation indexes and normalization values above and compare them against batch normalization. The above equations show that $\mathbb{E}[x_{i}]$ and $Var(x_{i})$ in layer normalization are the same for each component $x_{i}$ within a vector but are different for different vectors. This means that two variables are created in LayerNorm per vector. As LLMs will normalize way more than $2d$ vectors throughout the course of their training, the number of statistic variables used in LayerNorm is of the order of the number of tokens used to train the model. Thus, the information lost during layer normalization cannot be recovered by learning just $2d$ components of the $\bm{\alpha}$ and $\bm{\beta}$ learnable parameters. This shows that information once lost during layer normalization cannot be recovered during the scale-and-shift step, making the layer normalization process irreversible. In other words, the learnable parameters in LayerNorm cannot represent an identity transformation. To the best of our knowledge, the “irreversibility” of layer normalization has not been discussed previously in literature.
3.4 What’s so Special about the Uniform Vector?
We see that the definition of LayerNorm is innately linked with the uniform vector, possibly unintended by the original authors. Since the information along the uniform vector is being removed irreversibly during layer normalization (section 3.3), the layer normalization process implicitly assumes that the information along the uniform vector is either not important or that the model should not store information along that direction.
To understand what removing the component along the uniform vector does, we need to go back to the component-wise definition of LayerNorm. Removing the component along the uniform vector is the same as subtracting the mean of components of a vector from each component. If all components of a vector are equal, this operation reduces the vector to a vector of zeros. Thus, LayerNorm forces the model to learn hidden representations such that all components of the vector are not the same; otherwise, the vector is reduced to zero. A possible argument for enforcing this could be that if all components of a vector are the same, the vector loses expressive power. This is a possible conjecture offered by us to potentially justify such a choice. If this is indeed the explanation, the need for such a step should be justified by showing that the hidden representations within a language model have a propensity towards having all components equal to each other.
A different perspective to understand the implications of the “mean subtraction” step can be seen using the global picture of LayerNorm presented in this paper. Two randomly chosen vectors in high dimensional spaces are nearly orthogonal to each other with a very small spread. Since the “mean subtraction” step in LayerNorm means removing the component along the arbitrarily chosen uniform vector, the need for such a step would be justified if the hidden representations created by the model had significant components along this vector. The choice of removing information along the uniform vector seems like an unintended choice without much justification given in literature, including the original paper Ba et al. (2016). RMSNorm, a popular variant of LayerNorm used to train the Llama series of models, does not perform this step and performs just as well. This provides a strong motivation to explore the need for the “mean subtraction” step in LayerNorm (section 4.3).
Table 1: List of models used in experiments.
Model Name Model Dim (d) Num Params Num Layers Norm. Type
GPT-2 XL 1600 1.5B 48 Layer
GPT-Neo 1.3B 2048 1.3B 24 Layer
Pythia-1.4B 2048 1.4B 24 Layer
GPT-J 6B 4096 6.0B 28 Layer
Pythia-6.9B 4096 6.9B 32 Layer
Llama-2-7B 4096 7.0B 32 RMS
Llama-3-8B 4096 8.0B 32 RMS
4 Experiments
After a detailed theoretical analysis of the effects of layer normalization on a vector, we next study how LayerNorm affects the hidden representations of LLMs in practice. While prior works have studied the effects of LayerNorm on downstream model performance and training convergence (Ba et al., 2016; Xu et al., 2019; Zhang & Sennrich, 2019), there has been a surprising gap in the literature about studying the effect of LayerNorm on the internal representations of a model at inference time. To do so, we pass one million tokens from Wikipedia articles through various models and study how LayerNorm modifies a hidden vector in the representation space.
Methods and Models
We study the impact of LayerNorm on the hidden representations of 7 decoder-only LLMs across two size categories. The models used are listed in Table 1. GPT2XL, GPT-Neo-1.3B, and Pythia-1.4B represent the small LLM size category, whereas GPT-J-6B, Pythia-6.9B, Llama-2-7B, and Llama-3-8B represent the medium LLM size category. We pass one million tokens from Wikipedia articles through each model and capture the hidden representations for all tokens before and after each normalization layer. This is done for each layer inside the model, which creates many terabytes to be analyzed. Norms were calculated using the L2 norm, and angle differences in degrees were determined through cosine similarity calculations to measure the orientation changes induced by normalization.
As outlined in Table 1, we also use two types of normalization methods. Models like GPT-2 XL, GPT-Neo 1.3B, Pythia-1.4B, GPT-J 6B, and Pythia-6.9B employ LayerNorm (Ba et al., 2016), whereas Llama-2-7B and Llama-3-8B utilize RMSNorm (Zhang & Sennrich, 2019). The crucial difference between LayerNorm and RMSNorm is the subtraction of the mean vector (Figure 2(a)). In this paper, we study how LayerNorm and RMSNorm operate on the hidden representations.
(a) Residual Stream Pre-LN1 GPT-J
(b) Residual Stream Pre-LN1 Pythia 6.9
(c) Residual Stream Pre-LN1 Llama-3
(d) Residual Stream Post-LN1 GPT-J
(e) Residual Stream Post-LN1 Pythia 6.9
(f) Residual Stream Post-LN1 Llama-3
Figure 3: This figure shows the growing norm of the residual stream or hidden vectors at each layer (a-c) and how LayerNorm and RMSNorm regulate the growing norms (d-f) for GPTJ 6B, Pythia 6.9B and Llama3 8B.
4.1 Norm Stabilization
As a token representation passes through the different layers inside an LLM, the hidden representations accumulate due to residual connections (He et al., 2016). These hidden representations get added to the residual stream at each layer as shown in equation 6 and can cause the norm of the hidden vectors to grow. In fact, the norm of the hidden vectors at each layer grows disproportionately with large standard deviations, as can be seen in Figure 3 (a-c). The growing norm of the hidden vectors is extreme for some models more than others, as is seen for GPT-J and Pythia-6.9B. The growing norm of hidden representations for the remaining models from Table 1 and the effect of layer normalization can be seen in Figure 6 (appendix).
The significant role of LayerNorm and RMSNorm in stabilizing the growing norm of the hidden vectors at each layer can be seen in Figure 3 (d-f), which shows the norm of the hidden vectors after the first normalization step. This shows that normalization before sending vectors into the attention module significantly attentuates the vectors norms and reduces their variance. The standardization step in layer normalization first modifies the norm of each vector to $\sqrt{d}$, where $d$ is the dimensionality of the representation space. The norms of the standardized vectors are further modified marginally by the scale-and-shift steps. Layer normalization effects both the mean and the spread of the norms, where the standard deviations in norm reduce by factors between 10-100 depending on the model (Table 3 in appendix). These results show the crucial role that LayerNorm and RMSNorm play in stabilizing the intermediate hidden vectors in practice.
4.2 Rotation
While stabilizing the growing norm of hidden vectors is a major function of layer normalization, it also ends up rotating the hidden vectors. Figure 4 shows the angle between the original hidden vector and the post-normalization vector. As can be seen in Figure 4 (and in Figure 8 in appendix for the rest of the models), each layer is responsible for a fixed, non-trivial amount of rotation of hidden vectors in the representation space. Mean rotation angles in degrees between hidden and post-layer normalization vectors (LN1 and LN2) across all layers for various models are shown in Table 4 in the appendix. Thus, both LayerNorm and RMSNorm additionally rotate hidden vectors in representation space in addition to stabilizing the norms of the vectors.
4.3 LayerNorm versus RMSNorm : The need for “Mean Subtraction”
Apart from norm stabilization and rotation, a critical aspect of LayerNorm is its ability to orient the hidden vectors orthogonal to the uniform vector, as discussed in section 3. This is done by permanently (section 3.3) removing the component of every vector along the uniform vector and then normalizing the remaining vector. Such an action trains the model to operate orthogonal to the uniform vector, implying that the information along the uniform vector is not useful (section 3.4). RMSNorm, on the other hand, does no such thing and simply normalizes the existing vector, showing similar performance (Zhang & Sennrich, 2019).  Thus, a natural question to ask here is: “Is the mean subtraction step really needed?”
(a) GPT-J
(b) Pythia 6.9
(c) Llama-3
Figure 4: Rotation angle (in degrees) between the hidden vectors and Post-LN1 vectors across all layers for GPT-J, Pythia 6.9, Llama-3
(a) Hidden GPT-J
(b) Hidden Pythia 6.9
(c) Hidden Llama-3
(d) Post-LN1 GPT-J
(e) Post-LN1 Pythia 6.9
(f) Post-LN1 Llama-3
Figure 5: Distribution of angles (in degrees) between Hidden vectors (a-c) and post-normalization vectors (d-f) with the uniform vector for GPT-J, Pythia 6.9, Llama-3 for a layer picked at random (Layer 24).
The Llama model series (Touvron et al., 2023a; b) is the most prominent family of LLMs that use RMSNorm. While the fact that Llama models achieve state-of-the-art results across multiple measures (Dubey et al., 2024) shows that using RMSNorm instead of LayerNorm does not hurt performance, it is hard to isolate the role of the “mean subtraction” step in overall performance. In this paper, we answer this question mechanistically by finding justification for the need for such an operation. To do so, we find the angle between the hidden vector and the uniform vector just before and after layer normalization operations (equations 1 and 3). The results are shown in Figure 5 for GPT-J (6B), Pythia (6.9B), and Llama-3 (8B). We see that for all models in Figure 5, the angle between the hidden vectors and the uniform vectors is 90 degrees on average even before normalization, with a very small spread. This remains true post-normalization as well for all three models.
These results are surprising for two reasons. Firstly, the hidden representations for LayerNorm-based models (GPT-J and Pythia) are themselves orthogonal to the uniform vector even before the LayerNorm operation. This is true for 4 out of the 5 LayerNorm-based models studied in this paper except GPT2-XL (Figure 10 in the appendix). This shows that the “mean subtracting” operation of LayerNorm is redundant during inference. While indicating the redundancy of removing the component along the uniform vector during inference, it may be that the model learns to align the hidden representations orthogonal to the uniform vector during the training process, hence requiring the mean subtraction step. This is what makes the second observation even more surprising. We find that even for Llama-3 (8B) trained using RMSNorm, the hidden representations before normalization are also on average orthogonal to the uniform vector with a very small spread. In fact, they are more orthogonal to the uniform vector than for LayerNorm-based models, with a much smaller variance. The large variance for LayerNorm-based models is due to a long tail omitted in the figure. After RMSNorm, the hidden representations continue to be orthogonal to the uniform vector. This is also true for Llama-2 (7B) (Figure 10 in the appendix).
These results show that the “mean subtracting” operation in LayerNorm is redundant as the model learns to operate orthogonal to the uniform vector naturally during the training process, as seen in Llama-3 (8B) and Llama-2 (7B) trained with RMSNorm and for 4 out of 5 models trained using LayerNorm. The model naturally aligns its representations orthogonal to the uniform vector as expected in such high-dimensional spaces and, as a consequence, promotes diversity in hidden vector components. With this, we provide the first mechanistic evidence that the “mean subtraction” step in LayerNorm is dispensable, as we do not find hidden vectors having large components along the uniform vector. Moreover, the irreversibility of LayerNorm means that the information lost in the “mean subtraction” step is never recovered. Due to this, we advocate for using RMSNorm over LayerNorm, which is also computationally more efficient and leads to comparable downstream performance.
5 Conclusion
We present a detailed theoretical and empirical analysis of layer normalization on hidden vectors in representation space, with a focus on the global effects of the LayerNorm operation on a vector. We first show that the LayerNorm operation can be understood in three simple steps: removing the component of a vector along the uniform vector ($\bm{1}=[1,1,1,1,\cdots,1]^{T}\in\mathbb{R}^{d}$), normalizing the remaining vector, and scaling the resultant vector by $\sqrt{d}$. We then show that LayerNorm is an irreversible process—information along the uniform vector is removed during LayerNorm and cannot be recovered using the learnable parameters available in the formulation. This is in contrast with BatchNorm, where the network has the option of learning an identity operation. We then empirically show how LayerNorm regulates the norm and the orientation of hidden vectors during model inference. Finally, we show that the “mean subtraction” operation in LayerNorm is dispensable both during inference and for training, as shown by internal representations of RMSNorm-based Llama models.
Appendix A Appendix
This appendix includes detailed tables and figures that illustrate the effects of layer normalization on the hidden and post-layer normalization vectors across various models.
All experiments were conducted using NVIDIA A6000 GPUs with 48GB of GPU memory. This setup provided the necessary computational power to handle large-scale models and extensive data processing required for our study.
Plots in Figure 6 compare the norms of the residual streams before and after LN1 across the rest of the models not included in Figure 3, demonstrating the normalization’s effect. Plots in Figure 7, on the other hand, display the effect of LN2 on the residual stream norms across all models, highlighting the further stabilization achieved. Figure 8 shows the rotation angles between hidden vectors and post-LN1 vectors for the remaining models, illustrating the angular changes due to the first layer normalization. Figure 9 presents the rotation angles between hidden vectors and post-LN2 vectors for all models, and Figure 10 presents the angle frequency plots for the rest of the models.
Table 2: Average angles in degrees between a uniform vector and vectors from Hidden, Post-LN1, and Post-LN2 layers across all layers for each model
Model Name Hidden & Uniform (Mean $\pm$ Std) Post-LN1 & Uniform (Mean $\pm$ Std) Post-LN2 & Uniform (Mean $\pm$ Std)
GPT-2 XL 101.48 $\pm$ 3.20 89.95 $\pm$ 1.30 89.63 $\pm$ 1.30
GPT-Neo 1.3B 89.63 $\pm$ 1.84 88.85 $\pm$ 1.82 89.47 $\pm$ 1.84
Pythia-1.4B 89.88 $\pm$ 1.84 88.74 $\pm$ 1.81 89.39 $\pm$ 1.83
GPT-J 6B 88.74 $\pm$ 1.68 89.59 $\pm$ 1.69 -
Pythia-6.9B 89.27 $\pm$ 1.58 89.02 $\pm$ 1.57 88.97 $\pm$ 1.57
Llama-2-7B-hf 89.47 $\pm$ 0.15 89.48 $\pm$ 0.14 89.72 $\pm$ 0.15
Llama-3-8B 90.01 $\pm$ 0.15 90.15 $\pm$ 0.15 89.97 $\pm$ 0.15
Table 3: Average norms (L2) and standard deviations for hidden input vectors and post-layer normalization output vectors, alongside the maximum observed values for various models. ”Hidden” is an abbreviation for hidden input vectors.
Model Name Hidden (Mean $\pm$ Std) Max Hidden Post-LN1 (Mean $\pm$ Std) Max Post-LN1 Post-LN2 (Mean $\pm$ Std) Max Post-LN2
GPT-2 XL 236.85 $\pm$ 37.02 7186.44 21.42 $\pm$ 0.47 37.42 23.79 $\pm$ 0.49 38.51
GPT-Neo 1.3B 811.84 $\pm$ 40.29 5961.95 6.69 $\pm$ 0.26 32.99 13.79 $\pm$ 0.49 50.49
Pythia-1.4B 77.94 $\pm$ 13.17 1373.85 46.02 $\pm$ 1.11 81.54 35.68 $\pm$ 1.13 56.53
GPT-J 6B 110.88 $\pm$ 30.02 4757.86 49.44 $\pm$ 1.31 70.93 - -
Pythia-6.9B 216.04 $\pm$ 24.61 3119.49 79.65 $\pm$ 1.56 133.55 42.18 $\pm$ 1.03 65.39
Llama-2-7B-hf 44.30 $\pm$ 19.55 3287.86 26.21 $\pm$ 0.56 38.06 17.27 $\pm$ 0.42 30.76
Llama-3-8B 12.12 $\pm$ 3.15 549.19 29.46 $\pm$ 0.59 46.21 21.30 $\pm$ 0.49 39.67
Table 4: Average rotation angles in degrees between hidden and post-layer normalization vectors (LN1 and LN2) across all layers for various models, indicating the typical angular deviation introduced by normalization processes.
Model Name Hidden & Post-LN1 (Mean $\pm$ Std) Hidden & Post-LN2 (Mean $\pm$ Std)
GPT-2 XL 30.46 $\pm$ 1.68 32.68 $\pm$ 1.60
GPT-Neo 1.3B 49.09 $\pm$ 1.22 52.82 $\pm$ 1.34
Pythia-1.4B 12.35 $\pm$ 0.46 21.70 $\pm$ 0.90
GPT-J 6B 19.55 $\pm$ 0.98 -
Pythia-6.9B 11.02 $\pm$ 0.26 28.66 $\pm$ 0.58
Llama-2-7B-hf 17.13 $\pm$ 0.77 23.48 $\pm$ 1.16
Llama-3-8B 19.64 $\pm$ 0.62 24.55 $\pm$ 1.08
(a) Residual Stream Pre-LN1 GPT-2 XL
(b) Residual Stream Pre-LN1 GPT-Neo
(c) Residual Stream Pre-LN1 Pythia 1.4
(d) Residual Stream Pre-LN1 Llama-2
(e) Residual Stream Post-LN1 GPT-2 XL
(f) Residual Stream Post-LN1 GPT-Neo
(g) Residual Stream Post-LN1 Pythia 1.4
(h) Residual Stream Post-LN1 Llama-2
Figure 6: Pre and Post-LN1 plots of GPT-2 XL, GPT-Neo, Pythia 1.4 and Llama-2
(a) Residual Stream Post-LN2 GPT-Neo
(b) Residual Stream Post-LN2 Pythia 6.9
(c) Residual Stream Post-LN2 Llama-3
(d) Residual Stream Post-LN2 GPT-2 XL
(e) Residual Stream Post-LN2 Pythia 1.4
(f) Residual Stream Post-LN2 Llama-2
Figure 7: Affect of LN2 on stabilizing the norm in all models
(a) GPT-2 XL
(b) GPT-Neo
(c) Pythia 1.4
(d) Llama-2
Figure 8: Rotation angle (in degrees) between the hidden vectors and Post-LN1 vectors across all layers for GPT-2 XL, GPT-Neo, Pythia 1.4 and Llama-2
(a) Pythia 6.9
(b) Llama-3
(c) GPT-2 XL
(d) GPT-Neo
(e) Pythia 1.4
(f) Llama-2
Figure 9: Rotation angle (in degrees) between the hidden vectors and Post-LN2 vectors across all layers for all models
(a) Hidden GPT-2 XL
(b) Hidden GPT-Neo
(c) Hidden Pythia 1.4
(d) Hidden Llama-2
(e) Post-LN1 GPT-2 XL
(f) Post-LN1 GPT-Neo
(g) Post-LN1 Pythia 1.4
(h) Post-LN1 Llama-2
Figure 10: Distribution of angles (in degrees) between Hidden vectors (a-d) and post-normalization vectors (e-h) with the uniform vector for GPT-2 XL (Layer 43), GPT-Neo (Layer 18), Pythia 1.4 (Layer 18), and Llama-2 (Layer 24) for a layer picked at random
(a) Post-LN2 Pythia 6.9
(b) Post-LN2 Llama-3
(c) Post-LN2 GPT-2 XL
(d) Post-LN2 GPT-Neo
(e) Post-LN2 Pythia 1.4
(f) Post-LN2 Llama-2
Figure 11: Distribution of angles (in degrees) between (PostLN2 and Uniform) vectors for all models for a single layer chosen (Layer 43 for GPT-2 XL, Layer 18 for GPT-Neo and Pythia 1.4, and Layer 24 for Llama-2, Pythia 6.9, and Llama-3)